OpenStudy (anonymous):

lim as x-> 0+ of (sin(x+5pi/4))/(x)

sam (.sam.):

\[\lim_{x\to 0}\frac{\left( \sin \left(x+5 \frac{\pi }{4}\right)\right) }{x} \]